# SPICE: Smart Projection Interface for Cooking Enhancement

### v1

@@ 5. Results
Figure 5. Overview of the results obtained grouped by each one of the metrics considered. A) shows a boxplot with mean and standard deviations of the difficulty perceived for the experiment group that followed the recipe in the smartphone (orange), using SPICE (green), and the validation group (purple). B, C, D, and E show the same information for the Confidence, Taste, Efficiency, and Total Duration (in secs) metrics respectively.
Metric Experiment (smartphone) Experiment (SPICE) Validation
Efficiency (1-10) 7.631 8.526 6.300
Confidence (1-10) 5.310 7.895 5.400
Taste (1-10) 6.166 8.444 6.500
Difficulty (1-10) 7.894 8.000 7.800
Total duration (secs) 559.58 524.47 471.30
Number of stops 10.00 6.158 6.00
@@
Table 1. Experiment Results. Comparison of averages across Experiment (smartphone), Experiment (SPICE), and Validation groups for Efficiency, Confidence, Taste, and Difficulty, which are self-reported ratings by the participants, whereas the Total Duration and Number of stops are observed from the video recordings.
Table 1 and Fig. 5 show the comparative performance across the experiment (both on the smartphone and SPICE) and validation (only on SPICE) groups in several key metrics: Efficiency, and Total Duration, and Number of Stops. Efficiency, Confidence, Taste, and Difficulty were parameters assessed through a questionnaire, with participants self-rating themselves on a scale of 1 to 10. Duration was measured using the video footage.
Participants in the Experiment group (SPICE) reported higher Efficiency (8.526) and confidence (7.895) compared to both the Experiment (smartphone) and Validation groups (5.31 and 7.631, respectively). The Experiment group (SPICE) also rated the taste of dishes higher (8.444) compared to the Experiment (smartphone) group (6.166) and Validation group (6.500). The total duration of cooking tasks was less in the Experiment group (SPICE) (524.474 seconds) compared to the the Experiment group (smartphone) (559.588 seconds) and Validation group (471.300 seconds) indicating that the groups exposed to the SPICE interface, took less time to follow the recipe. Additionally, the number of stops decreased with the use of SPICE, as the Experiment (SPICE) and Validation groups showed fewer stops (6.158 and 6.00, respectively) compared to the Experiment (smartphone) group (10.00). These findings underscore the SPICE system’s potential in enhancing cooking experiences.
6. Discussion
The results from the comparative usability tests reveal significant insights into the effectiveness of the SPICE system. Participants in the Experiment group (SPICE) demonstrated higher efficiency and confidence levels, suggesting that the SPICE interface enhances users’ operational performance and self-assurance in cooking tasks. The improvement in taste scores could indicate a positive impact on the culinary outcomes. However, the participants who used SPICE also experienced a reduced total duration and fewer stops, highlighting the system’s ability to streamline the cooking process even though it requires two-handed interaction. This has strong implications for using TUIs and seamless integration of digital information in the immediate workspace without the use of a screen. Despite the initial positive findings, the difficulty ratings were similar across all groups, indicating that while the system offers benefits, it may be very novel and perceived as difficult to use, making the difficulty of completing the recipe higher. Further research with larger sample sizes would be necessary to generalize these findings and explore the nuances in user experience more comprehensively.
7. Future Work
@@
Assuming that the costs of installation (e.g., tracking system, communication network) are reduced, and the technology utilized in SPICE can be adapted to any kitchen environment, it is reasonable to believe that private kitchens, as well as professional culinary settings might incorporate a CPS to streamline and improve the process of cooking, by enhancing physical objects with their digital counterparts (e.g., data, instructions). It is possible to envision a future where a kitchen can adapt to its cook and be interacted with both digitally and physically simultaneously. Finally, more modalities such as gesture and voice recognition might be an interesting addition for the end user. The intention of the authors is to conduct further research in this direction.
8. Conclusion
@@
The intersection of the digital and the physical presents challenges in creating seamless, intuitive interfaces. Despite advances in Cyber-Physical Systems (CPS) and Human-Computer Interaction (HCI), a significant gap remains in effectively merging physical objects with digital interfaces. Through Tangible User Interfaces (TUIs), which give digital functionalities to physical objects, it becomes possible to enhance physical processes which usually require focus and two-handed interaction such as cooking. In this paper, we proposed SPICE (Smart Projection Interface for Cooking Enhancement); a prototype that provides a TUI to interact with and manipulate whilst cooking. Experiments conducted with 30 subjects assessed the efficacy of SPICE compared to traditional recipe-following methods through a comparative usability test. Results indicate that user confidence, efficiency, and taste improved by using the proposed system. In addition, the total duration of the cooking process was reduced. SPICE allowed users to employ both hands actively in the cooking process while still benefiting from digital information projected in their visual working space. This study provides insights on how to blend of physical objects and digital information, which might have beneficial implications not just in cooking, but also in other high-stakes environments where physical activities and hands-on interaction predominate such as in manufacturing, medicine, and education.

### v2

@@ IV Experiment Setup
Figure 5: Overview of the results obtained grouped by each one of the metrics considered. A) shows a boxplot with mean and standard deviations of the difficulty perceived for the experiment group that followed the recipe in the smartphone (orange), using SPICE (green), and the validation group (purple). B, C, D, and E show the same information for the Confidence, Taste, Efficiency, and Total Duration (in secs) metrics respectively.
Metric Experiment (smartphone) - (A) Experiment (SPICE) Validation - (B) Percentage Difference (A-B)
Efficiency (1-10) 7.631 8.526 6.300 -17.44%
Confidence (1-10) 5.310 7.895 5.400 +1.69%
Taste (1-10) 6.166 8.444 6.500 +5.41%
Difficulty (1-10) 7.894 8.000 7.800 -1.19%
Total duration (secs) 559.58 524.47 471.30 -15,77%
Number of stops 10.00 6.158 6.00 -40%
@@
TABLE I: Experiment Results. Comparison of averages across Experiment (smartphone), Experiment (SPICE), and Validation groups for Efficiency, Confidence, Taste, and Difficulty, which are self-reported ratings by the participants, whereas the Total Duration and Number of stops are observed from the video recordings. The last column depicts the percentage difference between the Experiment (smartphone) (A) and the Validation groups (B).
In the two types of tests, the recipe methods were evaluated according to the following metrics reported by the participants (on a scale from 1 to 10): Difficulty, a self-reported rating of recipe difficulty; Confidence, a self-reported rating of confidence in own performance in recipe execution; Taste, a self-reported rating of the final meal’s taste; Efficiency, a self-reported rating of efficiency in recipe execution. The user experience of the participants was also quantified using the following two metrics, recorded by the experimenters using video footage: Total Duration, the recorded time taken to complete the recipe in seconds, and Number of Stops, the recorded number of times a participant stopped cooking to look at the recipe.
The two types of tests were conducted on a 5-step recipe for making guacamole, using four basic ingredients (tomato, avocado, lemon, onion). Participants were provided with a cutting board, a knife, a fork, plates and napkins. Every recipe step required two-handed actions. All participants were video recorded while performing their cooking. After the recipe was completed, all participants filled out a self-evaluating questionnaire [^6] All questionnaires are available on the GitHub repository of the project. which tested for the variables of interest using 10-point Likert scales.
As shown in Fig. 4, 30 participants (ages 20 to 50) were divided into two groups: the Experiment group $(n=20)$ and the Validation group $(n=10)$. The Experiment group followed the guacamole recipe first by using a text-based recipe on a smartphone, then, after a two-week interval, by using the SPICE interface. The Validation group followed the recipe using only the SPICE interface. The Experiment group was composed of 7 women and 13 men, and the Validation group of 1 woman and 9 men. The self-rated level of familiarity with digital technology, which we obtained through participants’ responses to the questionnaire (on a scale of 1 to 10, from ”not familiar at all” to ”extremely familiar”), had an average of 6.17 and a mode of 8 for the Experiment group, and an average of 5.4 and a mode of 8 for the Validation group. Participants were mostly 4th-year undergraduate students, with a few staff members of the university also participating.
Both parametric (Paired Sample T-Test) and non-parametric (Wilcoxon Signed-Rank test) methods were considered for data analysis. The Shapiro-Wilk test was used to assess normality of the data distribution. If non-normality was detected $\alpha<0.05$, the Wilcoxon Signed-Rank test was chosen due to its suitability for small sample sizes and repeated measures.
V Results
Fig. 5 and Table I present the results of both experiment types: comparative usability test (both on the smartphone and SPICE) and validation test (only on SPICE).
Participants in the Experiment group reported higher Efficiency (mean: 8.526) and Confidence (mean: 7.895) when using SPICE compared to when using a smartphone (mean Efficiency: 7.631; mean Confidence: 5.31), and compared to the Validation group using SPICE (mean Efficiency: 6.3; mean Confidence: 5.4). The Experiment group also rated the Taste higher (mean: 8.444) when using SPICE compared to using the smartphone (mean: 6.166) and compared to the Validation group using SPICE (mean: 6.5).
The Duration of cooking tasks was lower in the Experiment group when using SPICE (mean: 524.47 seconds) than when using the smartphone (mean: 559.58 seconds). However, the total duration was lowest in the Validation group using SPICE (mean: 471.3 seconds).
Similar results were observed for the number of stops: the Experiment group when using SPICE had fewer stops (mean: 6.158) than when using the smartphone (mean: 10), and the Validation group using SPICE had the fewest stops (mean: 6), although only slightly lower than the Experiment group when using SPICE.
In terms of percentage difference between the Experiment group using the smartphone and the Validation group that was only exposed to SPICE (Table I), there was a decrease of 40% in the Number of Stops and a decrease of 15.77% in the Duration of cooking. Perceived Difficulty was reduced by 1.19% while self-perceived Taste and Confidence increased by 5.41% and 1.69% respectively. Finally, the self-perceived Efficiency decreased by 17.44%.
VI Discussion
Compared to using a text-based recipe on a smartphone, SPICE substantially increased the efficiency (lower duration, fewer stops) of cooking from a recipe for first-time users in our tests, both when using SPICE after using a smartphone for the same recipe, and when using SPICE without any prior introduction to the recipe. However, first-time users did not generally feel the efficiency improvement in their subjective experience, although they were more likely to report a feeling of greater efficiency if they used SPICE after having used a smartphone.
In the comparison between using a smartphone versus using SPICE, both without any prior introduction to the recipe, SPICE resulted in a shorter average cooking duration (15.77% reduction) and fewer average stops (40% reduction), highlighting the system’s ability to streamline the cooking process even though it requires two-handed interaction. This has strong implications for using TUIs and seamless integration of digital information in the immediate workspace without the use of a screen.
Participants that used SPICE without any prior introduction to the recipe self-reported much lower perceived efficiency than the smartphone users, despite their far shorter cooking time and fewer stops. This suggests that the unfamiliarity of the SPICE system for first-time users made them feel inefficient, despite their better performance. Studies with repeated usage would be needed to confirm whether user feelings of inefficiency were due only to unfamiliarity.
Although participants that used SPICE after a smartphone self-reported higher confidence, perceived efficiency, and perceived taste than when using the smartphone, this could be because participants were exposed to the same recipe twice, rather than because of users’ comparisons of the two methods. Further research with larger sample sizes, different recipes, and different exposure combinations would be necessary to explore nuances in user experience more comprehensively.
VII Future Work
@@
Assuming that the costs of installation (e.g., tracking system, communication network) are reduced, and the technology utilized in SPICE can be adapted to any kitchen environment, it is reasonable to believe that private kitchens, as well as professional culinary settings might incorporate a CPS to streamline and improve the process of cooking, by enhancing physical objects with their digital counterparts (e.g., data, instructions). It is possible to envision a future where a kitchen can adapt to its cook and be interacted with both digitally and physically simultaneously. More modalities such as gesture and voice recognition might be an interesting addition for the end user. Finally, more advanced CV techniques to detect and deal with uncountable ingredients (e.g., flour, salt, sugar) are natural extensions of this work. The intention of the authors is to conduct further research in this direction.
VIII Conclusion
@@
The intersection of the digital and the physical presents challenges in creating seamless, intuitive interfaces. Through Tangible User Interfaces (TUIs), which give digital functionalities to physical objects, it becomes possible to enhance physical processes which usually require focus and two-handed interaction such as cooking. In this paper, we proposed SPICE (Smart Projection Interface for Cooking Enhancement); a prototype that provides a TUI to interact with and manipulate whilst cooking. Experiments conducted with 30 subjects assessed the efficacy of SPICE compared to traditional recipe-following methods through a comparative usability test. Results indicate that the total duration and the number of stops during the cooking process was reduced. In addition, the user confidence, taste, and difficulty metrics also improved by using the proposed system. SPICE allowed users to employ both hands actively in the cooking process while still benefiting from digital information projected in their visual working space. This study provides insights on how to blend of physical objects and digital information, which might have beneficial implications not just in cooking, but also in other high-stakes environments where physical activities and hands-on interaction predominate such as in manufacturing, medicine, and education.
IX Acknowledgments
Project supported by a 2024 Leonardo Grant for Scientific Research and Cultural Creation from the BBVA Foundation. The BBVA Foundation accepts no responsibility for the opinions, statements and contents included in the project and/or the results thereof, which are entirely the responsibility of the authors.